### v4

@@ 3.2.2 Recomputing Attention for SnapKV
In implementing SnapKV, we generally follow the original procedure outlined in Section 2.4, except that eviction candidates in SnapStream exclude sink tokens. We define $Q_{\text{obs}}=[q_{L-L_{\text{obs}}},...,q_{L}]$ as the query projections of the most recent $L_{\text{obs}}$ tokens and $K_{\text{evict}}=[k_{L_{\text{sink}}+1},...,k_{L-L_{\text{recent}}}]$ as the key projections of the set of eviction candidates. Then, the top-$K$ indices $I$ are computed exactly as in Equations 1 and 2, except that $W\in\mathbb{R}^{B\times N\times L_{\text{obs}}\times L_{\text{evict}}}$ is defined as
@@
Once the sink tokens and the recent token ring buffer are constructed, the Top-$K$ KVs are gathered from the removed tokens using SnapKV and are appended at the end. Since the input sequence length $L$ may be less than $L_{sr}$, the Top-$K$ tokens may consist of some number of padding tokens $0\leq p\leq K$. We pass the number of padding tokens for each sequence as additional metadata from the prefill stage to decoding as an additional tensor valid_topk to be used for the construction of an attention mask that masks out any padding tokens in the Top-$K$ section of the cache during decoding.
@@ 4 Implementation
Our goal in formulating SnapStream is to bridge the implementation gap between KV cache compression and LLM deployments, particularly for frameworks that use static graphs. We demonstrate the effectiveness of our approach by applying SnapStream to a real deployment of a SoTA reasoning LLM, DeepSeek-R1-0528 deepseekr1, on the SN40L accelerator. In this section, we describe the kernel fusion and tensor sharding decisions, known as a “mapping”, for a deployment of DeepSeek-R1-0528 on a single node of $16\timesSN$40Ls. Separate mappings are created for prefill and decoding that optimize for each phase’s respective SLOs. We also show how SnapStream integrates seamlessly into these mappings.
@@ 4.1 Prefill
The minimum prefill deployment unit consists of a single node (16 RDUs), executed under 16-way tensor parallelism (TP16). Prefill runs at batch size 1 in a continuous batching style.
Figure 5 shows a high-level block diagram of the MoE prefill graph, including operators for SnapStream compression. The prefill graph is split into multiple fused kernels, with green boxes indicating kernel boundaries. The use of dataflow, large on-chip SRAM buffers, and a programmable interconnect enables unprecedented levels of operator fusion and pipelining. This results in increased operational intensity and reduced HBM traffic.
@@
  • StreamingLLM Ring Buffer: this kernel fuses four gather operations for Range 1 and Range 2 from Figure 4 for each of the key and value caches, as well as indexing for the sink tensors. All operations are equally sharded across 16 sockets along the head dimension $N$.
  • SnapKV Compression: this kernel fuses several operations described below. All operations are sharded along the head dimension $N$.
@@
Following the QKV and SnapStream kernels, the MHA kernel implements multi-head attention using the QKV projections of the previous sections. Similarly to FlashAttention, we fuse the entire $QK^{\intercal}$ multiplication, softmax and $PV$ multiplication into a single kernel to avoid materializing the full attention matrix $P$. We distribute all operators in the MHA kernel along the head dimension $N$ equally across 16 sockets, and tile $K$ and $V$ into blocks of 8192 within each socket.
@@
The entire FFN - including the shared and routed expert operators as well as the activation function - is fused into a single kernel. This allows all dynamically selected experts to execute without synchronization overhead at kernel boundaries and to deliver substantial performance improvements. The fused implementation also enables preloading the weights of the next expert while processing the current one, further reducing the Time to First Token (TTFT) in prefill.
@@
The kernel boundaries for the dense layers are similar to those in the MoE layers, except that in dense layers, the Output Proj and FFN are fused into a single kernel.
@@ 4.2 Decoding
The decoding stage runs on a single $16\timesSN$40L node. KV cache reuse significantly reduces the amount of computation required for decoding, effectively reducing the sequence dimension of most operations to 1. This allows us to compile even larger fused kernels for this stage, such that the entire computation graph is contained within two kernels: (a) a Multi-Head Latent Attention (MLA) kernel and (b) a Feed-Forward Network (FFN) kernel, as illustrated in Figure 7.
The MLA kernel includes QKV projections, attention computation, output projection, and router GEMM. The FFN section contains the shared and routed expert GEMMs. Both sections are implemented using fusion and spatial dataflow, as shown in Figure 6.
@@
Adding SnapStream during to the decoding graph is relatively simple, since all we need to change is the index used to do the KV cache update. This is done using the index in Equation 4.
@@ 5.1 Accuracy
We further evaluate the efficacy of SnapStream in a real deployment scenario with the mapping of DeepSeek-R1 described in Section 4 on $16\timesSN$40L sockets. We extend the SnapStream maximum KV cache length to 32k and evaluate the deployment on an improved long sequence benchmark, LongBench-v2, and two short-input reasoning tasks, AIME24 and LiveCodeBench. The results in Table 2 show that this real-life SnapStream deployment is able to maintain accuracy with respect to the full KV cache baselines, and in particular, SnapStream is able to outperform a similar deployment of fixed 32k sequence length without SnapStream on LongBench-v2.
@@ 5.2 Performance
Table 5 also shows the impact of KV-cache compression on decoding throughput when using the maximum batch size allowed for each sequence length. Throughput was measured on a single $16\timesSN$40L node and averaged over 10 runs. The reported values are scaled by 2.4 to account for the average number of tokens accepted per forward pass, assuming an 80% acceptance rate in Multi-Token Prediction (MTP). During decoding, all computations operate directly on the compressed KV-cache, eliminating the need for decompression and preserving compute efficiency. The combination of a smaller memory footprint and higher batch-level parallelism yields substantial end-to-end throughput gains — up to $4.3\times$ — for the evaluated configurations. These results highlight that SnapStream not only enables longer sequences and larger batches within the same memory budget but also directly improves decode performance through better hardware utilization and reduced memory bandwidth pressure.
@@ 6 Related Work
Sparse attention has received sustained research interest since the introduction of the Transformer architecture vaswani2023attentionneed. Early works are primarily motivated by improving the $O(L^{2})$ complexity of attention with sparse factorizations child2019generatinglongsequencessparse, fixed sparse patterns beltagy2020longformerlongdocumenttransformer, or locality-sensitive hashing kitaev2020reformerefficienttransformer. However, such methods require training from scratch with these modifications, leading to such methods falling out of favor as inroads were made by simply scaling pretraining tokens and parameter counts. Recent SoTA models have found some success by applying sliding-window attention to a subset of their layers gemmateam2024gemma2improvingopen; openai2025gptoss120bgptoss20bmodel.
@@
In response to the increased cost of LLM training, researchers have developed a variety of training-free KV cache compression methods. StreamingLLM xiao2024efficientstreaminglanguagemodels was originally developed as a method to enable LLMs to generate tokens beyond the maximum sequence length encountered during training, and allows for coherent text generation with a fixed KV cache budget, although accuracy may suffer. Quest tang2024questqueryawaresparsityefficient extends PagedAttention kwon2023efficientmemorymanagementlarge by offloading KV pages with low attention scores based on pre-computed statistics. InfLLM xiao2024infllmtrainingfreelongcontextextrapolation similarly offloads KV cache blocks to host memory and operates with a similar KV cache structure as SnapStream, but looks up Top-$K$ token blocks for every newly decoded token. SAGE-KV wang2025llmsknowdropselfattention and SnapKV li2024snapkvllmknowslooking perform one-time and length-thresholded KV cache compression, respectively, but do not distinguish between prefill and decoding. SnapStream builds on training-free compression literature by combining the one-time compression of SnapKV with efficient decoding using a fixed budget from StreamingLLM, fitting each technique into the prefill-decode scheduling of a production LLM server. We leave incorporating pre-computed, offloaded KV blocks/pages in the style of Quest and InfLLM to future work.
@@
However, training-free approaches invariably lead to some accuracy degradation, leading to a resurgence in trainable sparse attention. Cartridges eyuboglu2025cartridgeslightweightgeneralpurposelong propose compressing large corpuses into trainable KV caches with much smaller sequence lengths that may be optionally loaded using cached prefix lookups. Titans behrouz2024titanslearningmemorizetest propose to augment Transformers with a neural memory module that aims to memorize the data seen during training and improve efficiency and accuracy on long context tasks. DeepSeek-V2 introduced Multi-Head Latent Attention (MLA) deepseekai2024deepseekv2strongeconomicalefficient, which introduces shared low-rank projections of the query, key, and value vectors to reduce KV cache memory along the hidden dimension, while trading off for increased compute to project back into the full hidden dimension during decoding. Native Sparse Attention (NSA) yuan2025nativesparseattentionhardwarealigned proposes a similar KV cache structure to SnapStream, with hierarchical compression of past tokens, that is trainable and aims to reduce complexity for both prefilling and decoding during inference, but also the backward pass during training. DeepSeek Sparse Attention (DSA) deepseekv32exp proposes to finetune an existing dense attention model instead of training from scratch by training a token selection module, which still computes $O(L^{2})$ scores between tokens, but with fewer heads and smaller hidden dimensions than full attention, thus reducing the compute costs of prefill. While SnapStream already accommodates MLA by compressing the smaller, low-rank latent vectors along the sequence dimension instead of the full-rank KVs, admitting the other architectural modifications mentioned above requires additional modifications to fused kernels that we leave to future work.
@@ 7 Conclusion
In this paper, we present SnapStream, a training-free KV cache compression method that employs SnapKV compression during prefill and a StreamingLLM rolling KV cache window during decoding. We describe the details of a static graph implementation of this method, as well as the details for a production 16-socket mixed tensor-parallel and data-parallel mapping of DeepSeek-R1, on SN40Ls. This mapping uses SnapStream to reduce memory pressure and increase decoding throughput. We verify that SnapStream introduces minimal accuracy degradation on long sequence tasks for both Llama 3.1 8B Instruct and DeepSeek-R1, as well as additional reasoning and coding tasks for the latter model. Finally, we show how our SnapStream mapping enables us to achieve a $4.3\times$ improvement in decoding throughput, with at most a 5% increase in prefill latency. We hope this work inspires further study of KV cache compression methods in production LLM serving workloads and a variety of hardware accelerators, and leave a closer examination of hyperparameters and mappings of other sparse attention mechanisms to future work.
@@ Appendix A SnapStream Static Graph Pseudo-code
In this section, we show and describe PyTorch-style pseudocode for the operation of a SnapStream KV cache during prefill and decoding. We compare a naïve implementation of SnapStream with a rolling recent window buffer, and our instantiation with a ring buffer.
@@ A.1 Rolling Recent Window
Listing 1 shows an implementation of the KV cache compression process for a naïve implementation of SnapStream with a rolling window buffer, as in StreamingLLM. It uses the prefill_snapstream_topk method described in Listing 6 to extract the top-$K$ evicted tokens between the sink and recent tokens. Construction of the compressed KV cache is then a straightforward concatenation in Lines 25-26.
@@
Listing 2 shows an implementation of the decoding phase KV cache update using the naïve implementation of SnapStream with a rolling recent window buffer. In the case where cache_position (sequence length $L$ in previous sections) is less than max_kv_length, the update is a straightforward scatter (Lines 19-22). In the other case (Lines 24-32), we construct a new KV cache by concatenating slices of the sink, Top-$K$, and recent KVs. To accommodate different cache_position values at different batch indices in a static graph, we perform both KV cache updates and select between the two depending on the batch-indexed cache position on Line 34. In practice, the implementation of case 2 in this code snippet leads to materialization of many intermediate buffers that take up extra on-chip memory. For dataflow architectures in particular, this can lead to inefficient allocation of pattern memory units during placing and routing that prolong program load and increase latency.
@@ A.2 Ring Buffer
To avoid these issues, we implement the KV cache as a ring buffer instead, as described in Section 3.2. Listing 3 show torch-style pseudocode implementing the construction of such a ring buffer during SnapStream prefill.
@@
13 range1_lhb_factor = (input_length - sink_token_length +
@@
15 range1_lhb = (range1_lhb_factor * recent_token_length +
@@
18 range2_lhb = range1_lhb - recent_token_length
19 range2_lhb = max(range2_lhb, sink_token_length)
@@
21 select_offset = input_length - range1_lhb
@@
24 range2_indices = arange(start=range2_lhb,
25 end=range2_lhb + recent_token_length)[None, None, :]
@@
27 range2_k = gather(key_states, range2_indices, dim=2)
@@
29 range2_v = gather(value_states, range2_indices, dim=2)
@@
Then, we use prefill_streamingllm_slicing to construct the ring buffer for the full SnapStream KV Cache in Listing 4. This greatly simplifies the decoding stage to a single scatter for each of the key and value tensors in Listing 5. Note that in the pseudocode, Range 1 and Range 2 are swapped when compared to Figure 4.

### v5

@@ 3.2.2 Recomputing Attention for SnapKV
In implementing SnapKV, we generally follow the original procedure outlined in Section 2.4, except that eviction candidates in SnapStream exclude sink tokens. We define $Q_{\text{obs}}=[q_{L-L_{\text{obs}}},...,q_{L}]$ as the query projections of the most recent $L_{\text{obs}}$ tokens and $K_{\text{evict}}=[k_{L_{\text{sink}}+1},...,k_{L-L_{\text{recent}}}]$ as the key projections of the set of eviction candidates. Then, the top-$K$ indices $I$ are computed exactly as in Equations 1 and 2, except that $W\in\mathbb{R}^{B\times H\times L_{\text{obs}}\times L_{\text{evict}}}$ is defined as
@@
Once the sink tokens and the recent token ring buffer are constructed, the Top-$K$ KVs are gathered from the removed tokens using SnapKV and are appended at the end. Since the input sequence length $L$ may be less than $L_{sr}$, the Top-$K$ tokens may consist of some number of padding tokens $0\leq p\leq K$. We pass the number of padding tokens for each sequence as additional metadata from the prefill stage to decoding to be used for the construction of an attention mask that masks out any padding tokens in the Top-$K$ section of the SnapStream KV cache.
@@ 4 Implementation
Our goal in formulating SnapStream is to bridge the implementation gap between KV cache compression and LLM deployments, particularly for frameworks that use static graphs. We demonstrate the effectiveness of our approach by applying SnapStream to a real deployment of a state-of-the-art (SoTA) reasoning LLM, DeepSeek-R1-0528 deepseekr1, on the SN40L accelerator. In this section, we describe the kernel fusion and tensor sharding decisions, known as a “mapping”, for a single decoder layer of DeepSeek-R1-0528 on a single SN40L-16 node. Separate mappings are created for prefill and decoding that optimize for each phase’s respective SLOs. We also show how SnapStream integrates seamlessly into these mappings.
@@ 4.1 Prefill
The minimum prefill deployment unit consists of a single SN40L-16 node executing with 16-way tensor parallelism (TP16). Prefill runs at batch size 1 with continuous batching. DeepSeek-R1-0528 consists of two types of decoder layers: dense FFN and sparse Mixture-of-Experts (MoE) layers. Figure 5 shows a high-level block diagram of the MoE prefill graph, including operators for SnapStream compression. The prefill graph is split into multiple fused kernels, with green boxes indicating kernel boundaries. The use of dataflow, large on-chip SRAM buffers, and a programmable interconnect enables unprecedented levels of operator fusion and pipelining. This results in increased operational intensity and reduced HBM traffic.
@@
  • StreamingLLM Ring Buffer: this kernel fuses four gather operations for Range 1 and Range 2 from Figure 4 for each of the key and value caches, as well as indexing for the sink tensors. All operations are equally sharded across the number attention heads $H$ into 16 sockets.
  • SnapKV Compression: this kernel fuses several operations described below. All operations are sharded across the number of attention heads $H$. The operations are:
@@
Following the QKV, StreamingLLM Ring Buffer and SnapKV compression kernels, the MHA kernel implements multi-head attention using the QKV projections of the previous sections. Similarly to FlashAttention, we fuse the entire $QK^{\intercal}$ multiplication, softmax and $PV$ multiplication into a single kernel to avoid materializing the full attention matrix $P$. We distribute all operators in the MHA kernel along the number of attention heads $H$ equally across 16 sockets, and tile $K$ and $V$ into blocks of 8192 within each socket.
@@
The entire MoE FFN - including the shared and routed expert operators as well as the activation function - is fused into a single kernel. This allows all dynamically selected experts to execute without synchronization overhead at kernel boundaries and to deliver substantial performance improvements. The fused implementation also enables preloading the weights of the next expert while processing the current one, further reducing the Time to First Token (TTFT) in prefill.
@@
While we only describe the mapping for MoE layers here, the kernel boundaries for the dense layers are largely the same except that the Output Proj and FFN operator groups are fused into a single kernel.
@@ 4.2 Decoding
The decoding stage runs on a single SN40L-16 node. KV cache reuse significantly reduces the amount of computation required for decoding, effectively reducing the sequence dimension of most operations to 1. This allows us to compile even larger fused kernels for this stage, such that the entire computation graph is contained within two kernels: (a) a Multi-Head Latent Attention (MLA) kernel and (b) a Feed-Forward Network (FFN) kernel, as illustrated in Figure 7.
The MLA kernel includes QKV projections, attention computation, output projection, and the MoE router GEMM. The FFN section contains the shared and routed expert GEMMs. Both sections are implemented using fusion and spatial dataflow, as shown in Figure 6.
@@
Accommodating a SnapStream KV cache in the decoding graph is relatively simple, since all we need to change is the index used to do the KV cache update. This index is exactly the value $L_{rb}$ defined in Equation 4, and its computation easily fits into the fused kernel of the MLA section.
@@ 5.1 Accuracy
We further evaluate the efficacy of SnapStream in a real deployment scenario with the mapping of DeepSeek-R1 described in Section 4 on a SN40L-16 node. We extend the SnapStream maximum KV cache length to 32k and evaluate the deployment on an improved long sequence benchmark, LongBench-v2, and two short-input reasoning tasks, AIME24 and LiveCodeBench. The results in Table 2 show that this real-life SnapStream deployment is able to maintain accuracy with respect to the full KV cache baselines, and in particular, SnapStream is able to outperform a similar deployment of fixed 32k sequence length without SnapStream on LongBench-v2.
@@ 5.2 Performance
Table 5 also shows the impact of KV-cache compression on decoding throughput when using the maximum batch size allowed for each sequence length. Throughput was measured on a single SN40L-16 node and averaged over 10 runs. The reported values are scaled by 2.4 to account for the average number of tokens accepted per forward pass, assuming an 80% acceptance rate in Multi-Token Prediction (MTP). The combination of a smaller memory footprint and higher batch-level parallelism yields substantial end-to-end throughput gains — up to $4.3\times$ — for the evaluated configurations. These results highlight that SnapStream not only enables longer sequences and larger batches within the same memory budget but also directly improves decoding performance through better hardware utilization and reduced memory bandwidth pressure.
@@ 6 Related Work
Sparse attention has received sustained research interest since the introduction of the Transformer architecture vaswani2023attentionneed. Early works are primarily motivated by improving the $O(L^{2})$ complexity of attention with sparse factorizations child2019generatinglongsequencessparse, fixed sparse patterns beltagy2020longformerlongdocumenttransformer, or locality-sensitive hashing kitaev2020reformerefficienttransformer. However, such methods require training from scratch with these modifications, leading to such methods falling out of favor as inroads to better model performance were made by simply scaling pretraining tokens and parameter counts. Recent SoTA models have found some success by applying sliding-window attention to a subset of their layers gemmateam2024gemma2improvingopen; openai2025gptoss120bgptoss20bmodel.
@@
In response to the increased cost of LLM training, researchers have developed a variety of training-free KV cache compression methods. StreamingLLM xiao2024efficientstreaminglanguagemodels was originally developed as a method to enable LLMs to generate tokens beyond the maximum sequence length encountered during training, and allows for coherent text generation with a fixed KV cache budget, although accuracy may suffer. Quest tang2024questqueryawaresparsityefficient extends PagedAttention kwon2023efficientmemorymanagementlarge by offloading KV pages with low attention scores based on pre-computed statistics. InfLLM xiao2024infllmtrainingfreelongcontextextrapolation similarly offloads KV cache blocks to host memory and operates with a similar KV cache structure as SnapStream, but looks up Top-$K$ token blocks for every newly decoded token. SAGE-KV wang2025llmsknowdropselfattention and SnapKV li2024snapkvllmknowslooking perform one-time and length-thresholded KV cache compression respectively, but do not distinguish between prefill and decoding. SnapStream builds on training-free compression literature by combining the one-time compression of SnapKV with efficient decoding using a fixed budget from StreamingLLM, fitting each technique into the prefill-decode scheduling of a production LLM server. We leave incorporating pre-computed, offloaded KV blocks/pages in the style of Quest and InfLLM to future work.
@@
However, training-free approaches invariably lead to some accuracy degradation, leading to a resurgence in trainable sparse attention. Cartridges eyuboglu2025cartridgeslightweightgeneralpurposelong propose compressing large corpuses into trainable KV caches with much smaller sequence lengths that may be optionally loaded using cached prefix lookups. Titans behrouz2024titanslearningmemorizetest propose to augment Transformers with a neural memory module that aims to memorize the data seen during training and improve efficiency and accuracy on long context tasks. DeepSeek-V2 introduced Multi-Head Latent Attention (MLA) deepseekai2024deepseekv2strongeconomicalefficient, which introduces shared low-rank projections of the query, key, and value vectors to reduce KV cache memory along the hidden dimension. However, doing so requires increased compute to project these compressed KVs back into the full hidden dimension during decoding. Native Sparse Attention (NSA) yuan2025nativesparseattentionhardwarealigned proposes a similar KV cache structure to SnapStream, with hierarchical compression of past tokens, that is trainable and aims to reduce complexity for both training and inference. DeepSeek Sparse Attention (DSA) deepseekv32exp proposes to finetune an existing dense attention model instead of training from scratch by training a token selection module, which still computes $O(L^{2})$ scores between tokens, but with fewer heads and smaller hidden dimensions than full attention, thus reducing the compute costs of prefill. While SnapStream already accommodates MLA by compressing the smaller, low-rank latent vectors along the sequence dimension instead of the full-rank KVs, admitting the other architectural modifications mentioned above requires additional modifications to fused kernels that we leave to future work.
@@ 7 Conclusion
In this paper, we present SnapStream, a training-free KV cache compression method that employs SnapKV compression during prefill and a StreamingLLM rolling KV cache window during decoding. We describe the details of a static graph implementation of this method, as well as the details for a production 16-socket mixed tensor-parallel and data-parallel mapping of DeepSeek-R1-0528, on SN40Ls. This mapping uses SnapStream to reduce memory pressure and increase decoding throughput. We verify that SnapStream introduces minimal accuracy degradation on long sequence tasks for both Llama 3.1 8B Instruct and DeepSeek-R1, as well as additional reasoning and coding tasks for the latter model. Finally, we show how our SnapStream mapping enables us to achieve a $4.3\times$ improvement in decoding throughput, with at most a 5% increase in prefill latency. We hope this work inspires further study of KV cache compression methods in production LLM serving workloads and a variety of hardware accelerators, and leave a closer examination of hyperparameters and mappings of other sparse attention mechanisms to future work.
@@ Appendix A SnapStream Static Graph Pseudo-code
In this section, we show and describe PyTorch-style pseudocode for the operation of a SnapStream KV cache during prefill and decoding. We compare a naïve implementation of SnapStream with our ring buffer instantiation.
@@ A.1 Rolling Recent Window
Listing 1 shows a naïve implementation of the SnapStream KV cache compression process with a rolling window buffer, similar to StreamingLLM. It uses the prefill_snapstream_topk method described in Listing 6 to extract the top-$K$ evicted tokens between the sink and recent tokens. Construction of the compressed KV cache is then a straightforward concatenation in Lines 25-26.
@@
Listing 2 shows the corresponding naïve implementation of the SnapStream decoding phase KV cache update. In the case where cache_position (sequence length $L$ in previous sections) is less than max_kv_length, the update is a straightforward scatter (Lines 19-22). In the other case (Lines 24-32), we construct a new KV cache by concatenating slices of the sink, Top-$K$, and recent KVs. To accommodate different cache_position values at different batch indices in a static graph, we perform both KV cache updates and select between the two depending on the batch-indexed cache position on Line 34.
[…]
@@ A.2 Ring Buffer
To avoid these issues, we implement the KV cache as a ring buffer instead, as described in Section 3.2. Listing 3 shows torch-style pseudocode implementing the construction of such a ring buffer during SnapStream prefill.
@@
13 range2_lhb_factor = (input_length - sink_token_length +
@@
15 range2_lhb = (range2_lhb_factor * recent_token_length +
@@
17 range2_lhb = max(range2_lhb, sink_token_length)
18 range1_lhb = range2_lhb - recent_token_length
@@
21 select_offset = input_length - range2_lhb
22 range2_indices = arange(start=range2_lhb,
23 end=range2_lhb + recent_token_length)[None, None, :]
@@
26 range2_k = gather(key_states, range2_indices, dim=2)
@@
28 range2_v = gather(value_states, range2_indices, dim=2)
@@
The code in Listing 3 is then used as a subroutine in Listing 4 to construct the ring buffer for the full SnapStream KV Cache during prefill. This greatly simplifies the decoding stage to a single scatter for each of the key and value tensors, as can be seen in Listing 5.